OpenStudy (anonymous):

A farmer wants to enclose a rectangular field along a river on three sides. If 4,400 feet of fencing is to be used, what dimensions will maximize the enclosed area?

OpenStudy (amistre64):

F(x,y,L) = xy-L(2x+y-4400) Fx = y-2L = 0; y = 2L Fy = x - L = 0; x = L FL = 2x +y -440 2L +2L -440 = 0 4L = 440 L = 110 x = 110 y = 220

OpenStudy (anonymous):

the the side opposite the river is half of the fencing, the other two sides use 1/4 of the fence each

OpenStudy (anonymous):

you can bet your bottom dollar so when you see this on a test go right to the answer. the amount of fence is a red herring. if there was no river make a square

OpenStudy (anonymous):

@amistre you have a function of 3 variables! wow

OpenStudy (amistre64):

I do :) L is spose to be lamda tho

OpenStudy (anonymous):

oh, you actually lost me entirely.

OpenStudy (amistre64):

I read the section on LaGrange multipliers the other night and have been dying to use it ;)

OpenStudy (amistre64):

\[f(x,y)\text{ ; function to max/min}\]\[g(x,y)=0\text{ ; constraint function}\] \[F(x,y,\lambda)=f(x,y)-\lambda\ g(x,y)\text{ ; partial derivatives to 0}\]

OpenStudy (anonymous):

put P = perimeter of fence, x = length of side opposite river so other two sides are \[\frac{P-x}{2}\] and area is \[A(x)=\frac{x(P-x)}{2}=\frac{P}{2}x -\frac{x^2}{2}\] a parabola facing down whose vertex is at \[\frac{-b}{2a}=\frac{P}{2}\] so opposite side is half the perimeter and other two are one -fourth

OpenStudy (amistre64):

it has something to do with lining up the gradients of each function and making them equal

OpenStudy (anonymous):

you gotta be kidding!

OpenStudy (anonymous):

simple algebra does this, not calculus, but i am impressed for sure!

OpenStudy (anonymous):

Satellite73- how much would that be in feet